A shoulder fillet has the dimensions D = 2.5”, d = 2.0”, and r = 0.06” (Kt = 2.552). The shaft is made from steel with an ultimate strength of 118 ksi with a machined surface finish. The shaft receives bending a bending moment that varies from +35 kip-in to -5 kip in. (a) Estimate the fatigue life of this shaft (b) What residual compressive stress would improve the life to 750,000. (Use the Smith-Watson-Topper approach to account for mean stress.) (c) How much should the life improve if the surface were ground instead of machined? (Compute the new N and compare to part (a).

could you please solve it using the numbers in the question, please don't post a past version of it as those are incorrect
Answer #1

x = 0.06 KA = 2.552 atened steel 6u = 118k Si prof per ____ a) fatigue life of the shaft. Mean Stress, 6m= 6max tbmin 6m = 35- ka = surface tinish factor for machined Surface finish I Ka = a (but) as 4.51_b=-0.265 Ka = 4.SI (118) ka = 1.27 - -0. 265I Se=0.2 48 59 = 14.48 Ksi from Equation bm & ba al but logo (0.9 but A -22026 Endurance test الجلد - كل - - log on. L - 0 1libe From = 750 x 103 smith watson - Topper approach ben - --- yo:75X10 Oy lo 102 103 104 105 106 107 Reversals to failure →